Ryan scored 16, 22 and 24 points in three basketball games. What is the least number of points he must score in the next game to average at least 20 points per game for the four games?
Answer by checkley77(12569)

(16+22+24+x)/4>=20
(62+x)/4>=20
62+x>=4*20
x>=80-62
x>=18 points are needed in the 4th game to ave. 20 for the 4 games.
Proof:
(16+22+24+18)/4>=20
80/4>=20
20>=20